OpenStudy (laurenmyers1098):

I need help with this algebra 2 question? It's in the comments.

OpenStudy (laurenmyers1098):

Give an example of a conjunction that is used in the real world. Then, give another example of a disjunction that is used in the real world. Explain clearly why each example is either a conjunction or a disjunction. Then describe what the graph and interval notation for each would look like.

OpenStudy (laurenmyers1098):

(Please don't tell me this is in the wrong subject because I will screenshot this exact question in my math course and post it if I must)

OpenStudy (perl):

a conjunction example: to vote you need to be 18+ and a citizen of the united states . (either condition alone is not enough, you need both)

OpenStudy (perl):

oh you want an algebra problem that uses conjunction

OpenStudy (laurenmyers1098):

I think that I understand the conjunction/disjunction part - just any real world example of something with and and or. But what is confusing me is how to describe the graph and interval notation.

OpenStudy (perl):

conjunctions come up when we solve abolsute value inequalities For example if I say "Suitable temperature to operate this machine is within 50 degrees of zero degrees celsius"

OpenStudy (anonymous):

Conjunction: They smiled and laughed (they did both things) Disjunction: They either smiled or laughed (one or other) Graph would be a Venn diagram with 2 circles. Where they overlap is a conjunction, the rest of the area covered by both circles are your disjunctions, outside the circles and inside the bounding square is the rest of language! Interval notation: conjunction<disjunction<language

OpenStudy (anonymous):

in other worlds A "logical disjunction" simply means : " or " so; "A ∨ B " is read as "A or B ". Such a disjunction is false if both A and B are false. So, the result would be true if one or more operands are true. It would only be false if both or all it's operands are false. A "logical conjunction" simply means "and" so; the symbol is an inverted "v" " A (inverted v) B " are read as " A and B " . The conjunction would only be true if and only if both operands are true, otherwise it would be false. okay: properties under the conjunctions are: 1. "Associativity": the order of operations does not matter as long as the sequence of the operands is not changed. Eventhough parenthesis are changed, operands will still act the same. ex: (5+1) +5 = 5+ (1+5) 2. "Commutativity": to the ability to change the order of something without changing the end result. ex: 3 + 5 = 5 + 3 for addition; 5*3 = 3*5 for multiplication 3. "Distributivity": uses the "dristibutive law" ex: 5( 2 + 3) = (5*2) + (5*3) 4. "Idempotence": multiple applications of that operation will yield the same result. unary operations: - the absolute value operation is a unary operation on the real numbers - the opposite operation (-x) on the real numbers - the power operations (squaring, cubing, etc) on the real numbers - the factorial operation on the real numbers - the trigonometric functions (sin x, cos x, tan x, cot x, csc x, sec x) on the real numbers - the natural logarithm (ln x) on the real numbers - the logarithm of base 10 (log x) on the real numbers - logical negation on truth values These are examples of unary operations, and could be subject to "Idempotence" 5. "Monotonic" - actually these are used mostly for calculus.. There's much more to explain, but it would be best if there would be an example for it? have a nice day!

OpenStudy (laurenmyers1098):

@ChiefArnav Those are copied answers from yahoo answers..they didn't help me at all with this problem which is why I'm here :)

OpenStudy (perl):

example of conjunction: "To compete at wrestling in lightweight division an athlete must weight at least 90 pounds and no more than 130 pounds. " I am guessing on those figures, we can look them up on wikipedia

OpenStudy (perl):

these two sentences can be represented algebraically by w>= 90 and w <= 130 , where w = weight in pounds

OpenStudy (laurenmyers1098):

So if I were to describe the graph of that, I would say there would be a closed dot on 90 and a closed dot on 130 and it would be filled in in between? Meaning everything including and between 90 and 130 is a solution?

OpenStudy (perl):

right

OpenStudy (perl):

w >= 90 [======> w <= 130 <============= ] where they overlap is the compound solution 90 <= w <= 130

OpenStudy (laurenmyers1098):

And for a disjunction, I could say (sticking with the wrestling example for the sake of this thread) to compete in junior wrestling, an athlete must be below/exactly 13 or to compete in senior an athlete must be above/exactly 14? So the graph would be a closed dot on 13 and all the numbers to the left and a closed dot on 14 and all the numbers to the right?

OpenStudy (perl):

I was thinking of this example. in a certain lake, anglers may keep fish that are less than 15 inches long and fish that are greater than 25 inches long

OpenStudy (perl):

let L be the length of the fish. this produces two inequalities L < 15 or L > 25

OpenStudy (perl):

this is an example of a disjunction

OpenStudy (perl):

you can be more detailed about the type of fish, walleyes lets say

OpenStudy (perl):

or maybe bass :)

OpenStudy (laurenmyers1098):

Ah okay, I think I get it now. Thank you! My main issue was how to describe the graphs and interval notations but I understand now :) @perl